An arithmetic sequence has its 5th term equal to 22 and its 15th term equal to 62. Find a recursive formula and a formula for the nth term. Label which is which.
a5 =22 and a15 =62 but I don't know how to go on from that point..
Thanks so much for the help:)

Answer by Boreal(15235)

start with an=a1+(n-1)d
a5=a1+4d=22
a15=a1+14d=62
Therefore, subtracting
-10d=-40
d=4
a5=22=a1+16
a1=6
6,10,14,18,22,a10 is 42, a15 is 62.
an=6+(n-1)*4